# Adapting OpenAI’s CLIP Model for Few-Shot Image Inspection in Manufacturing Quality Control: An Expository Case Study with Multiple Application Examples

### v1

@@ 1 Introduction
The statistical process monitoring school, grounded in control chart theory, offers two compelling advantages. First, these methods require only in-control data for implementation, circumventing the often insurmountable challenge of collecting a representative sample of defective products. Second, they provide theoretical guarantees through their ability to aggregate information temporally, offering insights into shift location, timing, and magnitude. However, the practical implementation of such methods faces significant challenges. They typically require extensive in-control samples, verification of process stability, and specific mathematical formulations closely tied to particular quality data model assumptions, often resulting in methods tailored to specific applications with uncertain generalizability. Furthermore, the lack of accessible software packages poses a general challenge. When applied to image and video image data, traditional approaches typically begin with image preprocessing to extract specific features, followed by control charting to assess process stability over time. Stability can also be evaluated using functional or profile data, representing the in-control dynamics of the video image within the region of interest Megahed et al. (22), Menafoglio et al. (23), Colosimo (7), Colosimo et al. (8). Spatio-temporal modeling has proven effective for identifying out-of-control states, although it introduces additional complexity in the mathematical formulation Yan et al. (32).
@@
To examine the utility of our approach, we present five diverse case studies: metallic pan surface inspection (18), 3D printing extrusion profile analysis, stochastic textured surface evaluation (4, 5), automotive assembly inspection (6), and microstructure image classification for metal additive manufacturing (31). These cases represent a broad spectrum of manufacturing quality control challenges, from well-defined defect patterns to subtle variations in surface texture. Through these examples, we demonstrate that CLIP-based few-shot learning can offer a practical middle ground between traditional statistical process monitoring and modern computational approaches. Our examples indicate that CLIP-based few-shot learning can provide a new paradigm for image-based quality control in manufacturing environments due to its predictive performance and implementation simplicity.
@@ 2.1 The CLIP Model Architecture
Table 1: Popular image encoder models used with the CLIP model and their characteristics
@@ 2.2 Mathematical Foundations
$s(x,y)=\mbox{cosSim}(f_{\theta}(x),g_{\phi}(y))=\frac{f_{\theta}(x)\cdot g_{\phi}(y)}{\|f_{\theta}(x)\|\|g_{\phi}(y)\|}.$ (3)
@@
$c^{*}=\operatorname*{arg\,max}_{c\in\mathcal{C}}\mbox{cosSim}(f_{\theta}(x),g_{\phi}(t_{c}))$ (4)
@@
$s_{c}(x)=\max_{k\in{1,\ldots,K}}\mbox{cosSim}(f_{\theta}(x),f_{\theta}(x_{k}^{c})).$ (5)
@@ 3.1 Implementation Framework
The CLIP embedding process transforms each preprocessed image into a high-dimensional feature vector. For instance, the ViT-L/14 model maps each $336\times 336$ pixel image into a 768-dimensional embedding space, while ViT-B/32 produces 512-dimensional embeddings from $224\times 224$ pixel inputs. This transformation occurs through CLIP’s patching mechanism, where images are divided into $p\times p$ patches ($p=14$ for ViT-L/14, $p=32$ for ViT-B/32) before being processed by the transformer architecture. Note that the patching idea is similar to the region of interest (ROI) idea in image monitoring applications; however, unlike typical implementations of ROIs (21, 12, 24), the patches are all equal in size and are non-overlapping.
@@ 4.1 Dataset Description
We used their dataset structure, which includes a learning set (10 nominal and 10 defective images) and a testing set (50 nominal and 50 defective images). This setup allows us to directly compare our CLIP-based approach with their original results while examining both zero-shot and few-shot learning capabilities. It should be noted that these images were provided directly by the first author of Megahed and Camelio (18) and were not available online before our analysis. This ensures that the images were not part of CLIP’s pre-training dataset, and that this dataset can be used for evaluating CLIP’s generalization capabilities on truly unseen manufacturing data.
@@ 4.2 Zero-Shot Classification Performance
Building upon our illustrative example in Section 2, we expanded our zero-shot classification experiment in two ways. First, we simplified the classification to a binary problem (nominal vs. defective) rather than the five categories used in the illustration. Second, instead of analyzing a single image, we evaluated CLIP’s performance on the 100 test images (50 nominal, 50 defective). The use of multiple test images enables us to compute comprehensive classification metrics rather than just individual prediction probabilities (as we did in our illustrative example).
@@ 4.3 Few-Shot Classification Performance
We then implemented our few-shot learning approach using the same 10 nominal and 10 defective images for learning as those used in Megahed and Camelio (18). Table 3 displays these results alongside the original benchmark results from Megahed and Camelio (18). Our approach shows substantial improvements across most metrics. Moreover, there is no metric where the original approach outperformed the proposed method in predictive performance.
@@
Metric Original Results CLIP Few-Shot Improvement
Accuracy 0.880 0.940 $+$0.060
Sensitivity (Recall) 0.760 0.880 $+$0.120
Specificity 1.000 1.000 0.000
Precision 1.000 1.000 0.000
F1-Score 0.857 0.936 $+$0.079
AUC --- - 0.999 --- -
The AUC values were not reported in Megahed and Camelio (18).
These results highlight several key findings. First, our CLIP-based few-shot approach demonstrated significant improvement over the original benchmark across all metrics. Our results show gains in sensitivity ($+$0.120) while maintaining perfect specificity and precision. Second, the significant improvement in classification accuracy from zero-shot (0.630) to few-shot (0.940) validates our hypothesis that few-shot learning can effectively bridge the domain gap between CLIP’s pre-training and manufacturing applications. Finally, our simple methodology achieves better results without the need for the custom image processing pipelines and feature engineering required by the original approach. Instead, we rely on the standard image processing done by the CLIP model.
@@ 5.3 Few-Shot Learning and the Impact of Learning Set Size
We first evaluated CLIP’s few-shot learning performance using a baseline set of 20 examples (10 nominal, 10 defective). This setup was selected since it is similar to what we did in the previous case study. Table 5 presents these initial results, where we can see that the predictive performance is poor. We suspect that this is mainly due to the low resolution images available in this case study; the extrusion profile images were only $85\times 85$ pixels, significantly smaller than CLIP’s expected input dimensions ($336\times 336$ for the ViT-L/14 encoder model). Note that throughout this case study, the number of nominal images was equal to the number of defective images (which was equally split between over and under extruded defects). Hereafter, we use the term “per class” to denote the nominal and overarching defective classes.
@@
Accuracy Sensitivity Specificity Precision F1-Score AUC
0.570 0.400 0.740 0.600 0.480 0.620
To investigate whether additional examples could improve these baseline results, we conducted a comprehensive analysis varying the learning set size from 10 to 350 nominal examples (i.e., total learning sample size is double these numbers) while maintaining a fixed test set for consistent comparison. Our results, depicted in Figure 6, revealed three key findings. First, even with just 10 examples per class, the model achieved an accuracy of 0.57 with 40% of the defects (sensitivity of 0.4) and 74% of the nominal cases (specificity of 0.74) correctly identified. Second, we observed improved performance between 10 and 100 per class examples for some classification metrics. For example, accuracy improved from 0.57 to 0.71, and AUC increased from 0.62 to 0.75; however, specificity remained relatively stable. Finally, performance improvements plateaued beyond 200 per class examples. The maximum values for our predictive metrics were achieved at 350 nominal learning examples, with an accuracy of 0.80, specificity of 0.81, and sensitivity of 0.78.
@@ 6.1 Dataset Description
Our third case study examines stochastic textured surfaces (STS), a unique manufacturing quality control data class that presents distinct challenges from traditional profile monitoring. STS data include woven textiles, surface metrology data from machined, cast, or formed metal components, and microscopy images of material microstructures (4). As noted by Bui and Apley (4), STS data present several challenges: (a) unlike our first case study of metallic surfaces where the gold standard is simply a non-textured surface of constant intensity or cases with deterministically repeated patterns where computer-aided design models can serve as references, STS has no definitive “gold standard” image; (b) under normal process conditions, there exist infinitely many valid images that exhibit identical statistical properties while being completely different at the pixel level; and (c) these images cannot be easily aligned, transformed, or warped into a standard reference image due to their inherent stochasticity. These challenges render both traditional profile monitoring and image inspection/monitoring approaches inapplicable (22, 4). This fundamental limitation of traditional approaches motivates our investigation of CLIP’s potential for STS image inspection, i.e., can vision transformers (with few-shot examples) help alleviate the aforementioned limitation of traditional approaches?
To systematically evaluate CLIP’s performance on STS image classification, we utilized the spc4sts package (5) to generate a comprehensive dataset of simulated textile images. This approach allowed for the controlled generation of a large-scale data set, including nominal and defective samples with known characteristics and precise manipulation of defect types (local vs. global) and magnitudes. Our experimental dataset included 1,000 nominal images representing standard weave patterns, 500 images with local defects characterized by localized disruptions in the weave pattern, and 500 images with global defects involving systematic shifts in weave parameters. Each image was generated at $250\times 250$ pixels following Bui and Apley (5)’s recommended parameters. Nominal images were generated using spatial autoregressive parameters $\phi_{1}=0.6$ and $\phi_{2}=0.35$. Local defects were imposed using the package’s defect generation function, and global defects were simulated by reducing both parameters by 5%. Figure 7 shows representative examples of the nominal and two defect categories. Note the subtle visual differences between the cases.
@@ 6.3 Model Selection Impact on Few-Shot Learning
We suspect that one contributor to the poor performance of the zero-shot approach is the complexity of the STS images. To investigate whether a more detailed vectorization of the images might improve performance, we compare two image encoders: ViT-L/14 and ViT-B/32. The ViT-L/14 encoder uses a $14\times 14$ pixel patch size, giving a more fine-tuned vectorization of the images than the ViT-B/32 which uses a $32\times 32$ pixel patch. Figure 8 shows how classification metrics evolved with increasing learning set size for both models. Note that throughout this case study, the number of nominal images was equal to the number of defective images (which was equally split between local and global defects).
@@
These findings highlight the importance of encoder model selection for STS inspection tasks. The superior performance of ViT-L/14 suggests that its larger model capacity and finer-grained patch size ($14\times 14$ vs $32\times 32$) allow for capturing the subtle patterns that characterize STS defects. This aligns with theoretical expectations, as smaller patch sizes should better preserve the local spatial correlations that define STS patterns. While the input images ($250\times 250$ pixels) were smaller than ViT-L/14’s designed input size ($336\times 336$ pixels), the model still performed well. This finding suggests that the model’s architecture is robust to moderate downscaling when the resolution remains sufficient to capture relevant texture patterns.

### v2

@@ 1 Introduction
The statistical process monitoring school, grounded in control chart theory, offers two compelling advantages. First, these methods require only in-control data for implementation, circumventing the often insurmountable challenge of collecting a representative sample of defective products. Second, they provide theoretical guarantees through their ability to aggregate information temporally, offering insights into shift location, timing, and magnitude. However, the practical implementation of such methods faces significant challenges. They typically require extensive in-control samples, verification of process stability, and specific mathematical formulations closely tied to particular quality data model assumptions, often resulting in methods tailored to specific applications with uncertain generalizability. Furthermore, the lack of accessible software packages poses a general challenge. When applied to image and video image data, traditional approaches typically begin with image preprocessing to extract specific features, followed by control charting to assess process stability over time. Stability can also be evaluated using functional or profile data, representing the in-control dynamics of the video image within the region of interest (27, 28, 8, 9). Spatio-temporal modeling has proven effective for identifying out-of-control states, although it introduces additional complexity in the mathematical formulation (38).
@@
To examine the utility of our approach, we present five diverse case studies: metallic pan surface inspection (23), 3D printing extrusion profile analysis, stochastic textured surface evaluation (5, 6), automotive assembly inspection (7), and microstructure image classification for metal additive manufacturing (37). These cases represent a broad spectrum of manufacturing quality control challenges, from well-defined defect patterns to subtle variations in surface texture. Through these examples, we demonstrate that CLIP-based few-shot learning can offer a practical middle ground between traditional statistical process monitoring and modern computational approaches. Our examples indicate that CLIP-based few-shot learning can provide a new paradigm for image-based quality control in manufacturing environments due to its predictive performance and implementation simplicity.
@@ 2.1 The CLIP Model Architecture
Table 1: Popular image encoder models used with the CLIP model and their characteristics.
@@ 2.2 Mathematical Foundations
$s(x,y)=\text{cosSim}\big{(}f_{\theta}(x),g_{\phi}(y)\big{)}=\frac{f_{\theta}(x)\cdot g_{\phi}(y)}{\|f_{\theta}(x)\|\|g_{\phi}(y)\|}.$ (3)
@@
$c^{*}=\operatorname*{arg\,max}_{c\in\mathcal{C}}\text{cosSim}\big{(}f_{\theta}(x),g_{\phi}(t_{c})\big{)}$ (4)
@@
$s_{c}(x)=\max_{k\in{1,\ldots,K}}\text{cosSim}\big{(}f_{\theta}(x),f_{\theta}(x_{k}^{c})\big{)}.$ (5)
@@ 3.1 Implementation Framework
The CLIP embedding process transforms each preprocessed image into a high-dimensional feature vector. For instance, the ViT-L/14 model maps each $336\times 336$ pixel image into a 768-dimensional embedding space, while ViT-B/32 produces 512-dimensional embeddings from $224\times 224$ pixel inputs. This transformation occurs through CLIP’s patching mechanism, where images are divided into $p\times p$ patches ($p=14$ for ViT-L/14, $p=32$ for ViT-B/32) before being processed by the transformer architecture. Note that the patching idea is similar to the region of interest (ROI) idea in image monitoring applications; however, unlike typical implementations of ROIs (26, 17, 29), the patches are all equal in size and are non-overlapping.
@@ 3.2 Rationale for the Selected Cases and Computational Experiment Overview
3.3 Selected Benchmarks for our CLIP Implementation and their Rationale
We compare our proposed CLIP-based few-shot learning approach against three references: (1) Published or Industry Benchmarks; (2) a MobileNetV2-based prototype classifier; and (3) Google Cloud Vertex AI’s AutoML Vision.
Published and Industry Benchmarks.
We report original metrics alongside our CLIP results when established baselines exist for a case study. These benchmarks measure progress relative to proven few-shot protocols and industrial best practices. This ensures meaningful and contextually relevant comparisons.
MobileNetV2 Few-Shot Prototype.
We use MobileNetV2 as a lightweight baseline for few-shot image classification (34). It is a well-established method in the literature (with over 29,000 Google Scholar citations) that preceded the release of the CLIP model. MobileNetV2 uses depthwise-separable convolutions, inverted residual blocks, and linear bottlenecks to reduce parameters and computational complexity while retaining representational capacity. The model is pretrained on ImageNet (10) and provides robust feature embeddings that transfer readily to downstream tasks.
We freeze the entire MobileNetV2 backbone and extract per-image embeddings via its features trunk. We compute class prototypes by averaging the embeddings of the nominal and defective training samples. Test images are assigned to the nearest prototype using cosine similarity. The distance to the defective prototype serves as a continuous anomaly score.
AutoML Vision.
We leverage AutoML Vision within Google Cloud’s Vertex AI to benchmark against a turnkey commercial solution. This assesses the value proposition of a fully automated, managed service. AutoML automates the model development process by exploring architectures such as MnasNet (derived from MobileNetV2), EfficientNet, and other CNN variants (12, 14). It applies automatic hyperparameter tuning to optimize performance for our task. AutoML incurs a cost of \$27.72 (USD) per case. Both the CLIP- and MobileNetV2-based prototypes require no paid computation.
We used the same test dataset for our CLIP based few-shot approach, MobileNetV2-based prototype classifier, and AutoML Vision to ensure fair evaluation. Vertex AI AutoML requires separate training, validation, and test datasets. We partitioned our original learning data by allocating 90% for training and 10% for validation. This follows Vertex AI’s documentation (13). The validation set tunes the model’s hyperparameters based on the model’s performance. Our few-shot CLIP-based method and the MobileNetV2 prototype classifier require only learning and test sets. They involve no model optimization or hyperparameter tuning during training.
AutoML Vision operates on Google-managed infrastructure designed for scalable and high-performance training. The exact hardware is not disclosed. All few-shot learning experiments were conducted in a lightweight computing environment. We used a Google Colab instance with a 2-thread AMD EPYC virtual CPU. This disparity ensures that observed differences in training runtime or computational cost are not due to hardware advantages on our method’s part.
@@ 4.1 Dataset Description
We adopted the original dataset split from the first author of Megahed and Camelio (23), using 10 nominal and 10 defective images for our few-shot learning. For the test set, we began with the 50 nominal and 50 defective images provided by Megahed and Camelio (23). Then, we removed training images, which yielded a test set of 39 nominal and 39 defective images. All images were never released publicly and hence, fall outside CLIP’s pre-training corpus.
@@ 4.2 Zero-Shot Classification Performance
Building upon our illustrative example in Section 2, we expanded our zero-shot classification experiment in two ways. First, we simplified the classification to a binary problem (nominal vs. defective) rather than the five categories used in the illustration. Second, instead of analyzing a single image, we evaluated CLIP’s performance on the 78 test images (39 nominal, 39 defective). The use of multiple test images enables us to compute comprehensive classification metrics rather than just individual prediction probabilities (as we did in our illustrative example).
@@ 4.3 Few-Shot Classification Performance
We then implemented our few-shot learning approach using the same 10 nominal and 10 defective images for learning as those used in Megahed and Camelio (23). Table 3 displays these results, which demonstrate that our CLIP-based few-shot pipeline delivers exceptionally strong performance with a learning set of size only 20. With an overall accuracy of 0.91 and an AUC of 0.998, coupled with perfect specificity and precision, our approach eliminates false positives while maintaining a high true-positive rate (sensitivity = 0.82) and balanced F1-score of 0.9. Compared to the zero-shot baseline, the few-shot adaptation boosts recall more than threefold and roughly doubles the F1-score. Our few-shot learning adaptation of CLIP illustrates how even a small set of learning images can bridge CLIP’s domain gap and meet or exceed typical industrial inspection standards.
@@
Method Accuracy Sensitivity Specificity Precision F1-score AUC
CLIP Few-Shot 0.910 0.821 1.000 1.000 0.901 0.998
4.4 Comparison with Other Benchmark Approaches
Table 4 compares our CLIP-based few-shot pipeline against three reference methods: the original benchmark from (23), a MobileNetV2-based few-shot prototype (MobileNetV2-FS), and Google Cloud Vertex AI’s AutoML Vision. Our approach achieves the highest accuracy (0.910). This improves on the original benchmark (0.880), MobileNetV2-FS (0.654), and AutoML (0.795). CLIP’s sensitivity more than doubles that of MobileNetV2-FS (0.308) and exceeds AutoML’s 0.641. All methods maintain near-perfect specificity and precision.
Our pipeline runs in 9.1 minutes at zero monetary cost. Most of this time is spent loading the CLIP model and conducting the learning process, rather than classification, which occurs in near real-time. That being said, the MobileNetV2-FS approach only required 1.4 minutes for the entire process of few-shot learning and classification. On the other hand, the AutoML was significantly slower (93.0 minutes) and more expensive (\$27.72). Note that the original benchmark of Megahed and Camelio (23) did not report AUC, runtime, or costs.
In this case study, our CLIP-based baseline outperforms both academic and commercial alternatives with minimal code and no fees. It preserves perfect false-positive control. Furthermore, it delivers practical runtime performance despite the model loading overhead.
Table 4: Performance and runtime comparison for metallic pan surface inspection. Bolded values denote the best-performing model for each metric.
Model Accuracy Sensitivity Specificity Precision F1 AUC Runtime (min)
CLIP Few-Shot 0.910 0.821 1.000 1.000 0.901 0.998 9.1
Benchmark* 0.880 0.760 1.000 1.000 0.857 - - - -
MobileNetV2-FS 0.654 0.308 1.000 1.000 0.471 0.620 1.4
AutoML 0.795 0.641 0.949 0.926 0.758 - - 93.0
*Results in (23) were based on 100 test images. The other approaches use 78 test images after removing any overlap with the training set. - - indicates the value is not available/reported.
@@ 5.3 Few-Shot Learning and the Impact of Learning Set Size
We first evaluated CLIP’s few-shot learning performance using a baseline set of 20 examples (10 nominal, 10 defective). This setup was selected since it is similar to what we did in the previous case study. Table 6 presents three initial results, where we can see that the predictive performance is poor. We suspect that this is mainly due to the low resolution images available in the case study; the extrusion profile images were only $85\times 85$ pixels, significantly smaller than CLIP’s expected input dimensions ($336\times 336$ for the ViT-L/14 encoder model). Note that throughout this case study, the number of nominal images was equal to the number of defective images (which was equally split between over- and under-extruded defects). Hereafter, we use the term “per class” to denote the nominal and overarching defective classes.
@@
Method Accuracy Sensitivity Specificity Precision F1-score AUC
CLIP Few-Shot 0.570 0.400 0.740 0.600 0.480 0.620
To investigate whether additional examples could improve the baseline results, we conducted a comprehensive analysis varying the learning set size from 10 to 350 nominal examples (i.e., total learning sample size is double these numbers) while maintaining a fixed test set for consistent comparison. Our results, depicted in Figure 6, revealed three key findings. First, even with just 10 examples per class, the model achieved an accuracy of 0.57 with 40% of the defects (sensitivity of 0.4) and 74% of the nominal cases (specificity of 0.74) correctly identified. Second, we observed improved performance between 10 and 100 per class examples for some classification metrics. For example, accuracy improved from 0.57 to 0.71, and AUC increased from 0.62 to 0.75; however, specificity remained relatively stable. Finally, performance improvements plateaued beyond 200 per class examples. The maximum values for our predictive metrics were achieved at 350 nominal learning examples, with an accuracy of 0.80, specificity of 0.81, and sensitivity of 0.78.
@@
5.4 Comparison with Other Benchmark Approaches
For this case study, no published results are available. We therefore compare only two methods against our approach under the same few-shot training regime (100 labeled images equally split among classes). AutoML is trained on 90 images, with 10 reserved for validation. The MobileNetV2-based few-shot prototype and our CLIP-based few-shot pipeline both use all 100 images. Table 7 captures the results of the comparison. MobileNetV2-FS attains the highest accuracy (0.805), specificity (0.890), precision (0.867), F1-score (0.787), and AUC (0783). MobileNetV2-FS leads in raw performance. CLIP-based method delivers reasonable sensitivity and discriminative power (AUC = 0.746) at zero cost and with minimal code, but it lags behind MobileNetV2-FS on overall accuracy, precision, and F1.
Table 7: Performance and runtime comparison for 3D printing extrusion profile inspection. Bolded values denote the best-performing model for each metric.
Model Accuracy Sensitivity Specificity Precision F1 AUC Runtime (min)
CLIP Few-Shot 0.700 0.710 0.690 0.696 0.703 0.746 17.0
MobileNetV2-FS 0.805 0.720 0.890 0.867 0.787 0.783 1.3
AutoML 0.775 0.800 0.750 0.762 0.781 - - 103.0
- - indicates the value is not available/reported.
@@ 6.1 Dataset Description
Our third case study examines stochastic textured surfaces (STS), a unique class of manufacturing quality control data that presents distinct challenges compared to traditional profile monitoring. STS data include woven textiles, surface metrology data from machined, cast, or formed metal components, and microscopy images of material microstructures (5). As noted by Bui and Apley (5), STS data present several challenges: (a) unlike our first case study of metallic surfaces where the gold standard is simply a non-textured surface of constant intensity or cases with deterministically repeated patterns where computer-aided design models can serve as references, STS has no definitive “gold standard” image; (b) under normal process conditions, there exist infinitely many valid images that exhibit identical statistical properties while being completely different at the pixel level; and (c) these images cannot be easily aligned, transformed, or warped into a standard reference image due to their inherent stochasticity. These challenges render both traditional profile monitoring and image inspection/monitoring approaches inapplicable (27, 5). This fundamental limitation of conventional approaches motivates our investigation of CLIP’s potential for STS image inspection, i.e., can vision transformers (with few-shot examples) help alleviate the aforementioned limitation of traditional approaches?
To systematically evaluate CLIP’s performance on STS image classification, we utilized the spc4sts package (6) to generate a comprehensive dataset of simulated textile images. This approach enabled the controlled generation of a large-scale dataset, including both nominal and defective samples with known characteristics, as well as the precise manipulation of defect types (local vs. global) and magnitudes. Our experimental dataset included 1,000 nominal images representing standard weave patterns, 500 images with local defects characterized by localized disruptions in the weave pattern, and 500 images with global defects involving systematic shifts in weave parameters. Each image was generated at $250\times 250$ pixels following Bui and Apley (6)’s recommended parameters. Nominal images were generated using spatial autoregressive parameters $\phi_{1}=0.6$ and $\phi_{2}=0.35$. Local defects were imposed using the package’s defect generation function, and global defects were simulated by reducing both parameters by 5%. Figure 7 shows representative examples of the nominal and two defect categories. Note the subtle visual differences between the cases.
@@ 6.3 Model Selection Impact on Few-Shot Learning
We suspect that one contributor to the poor performance of the zero-shot approach is the complexity of the STS images. To investigate whether a more detailed vectorization of the images might improve performance, we compare two image encoders: ViT-L/14 and ViT-B/32. The ViT-L/14 encoder uses a $14\times 14$ pixel patch size, giving a more fine-tuned vectorization of the images than the ViT-B/32, which uses a $32\times 32$ pixel patch. Figure 8 shows how classification metrics evolved with increasing learning set size for both models. Note that throughout this case study, the number of nominal images was equal to the number of defective images (which was equally split between local and global defects).
@@
These findings highlight the importance of encoder model selection for STS inspection tasks. The superior performance of ViT-L/14 suggests that its larger model capacity and finer-grained patch size ($14\times 14$ vs $32\times 32$) allow for capturing the subtle patterns that characterize STS defects. This finding aligns with theoretical expectations, as smaller patch sizes should better preserve the local spatial correlations that define STS patterns. While the input images ($250\times 250$ pixels) were smaller than ViT-L/14’s designed input size ($336\times 336$ pixels), the model still performed well. This finding suggests that the model’s architecture is robust to moderate downscaling when the resolution remains sufficient to capture relevant texture patterns.
@@
6.4 Comparison with Other Benchmark Approaches
In the third case study on STS inspection (Table 9), the original training set of 100 images was used in all three approaches. CLIP-based model achieves the highest overall accuracy (0.975), F1-score (0.975), specificity (0.990), and AUC (0.996). Furthermore, its sensitivity of 0.960 is quite competitive with the highest sensitivity achieved (0.970) using the AutoML approach. For this case study, our zero-cost approach offers the best trade-off between accuracy, runtime, and implementation simplicity compared to the two reference approaches.
As a contextual reference, Bui and Apley (5) utilized a control-chart based method for STS images; thus, they reported the average power (i.e., sensitivity) between 0.858 and 1.000 at a fixed Type I error of $\alpha=0.003$ (a controlled specificity of 0.997). While our method and theirs have utilized different test sets and evaluation criteria, these figures underscore that our CLIP-based approach attained sensitivity and specificity comparable to those of established statistical inspection techniques, which further support its viability as a lightweight, few-shot baseline for texture-based quality control.
Table 9: Performance and runtime comparison for stochastic textured surfaces. Bolded values denote the best-performing model for each metric.
Model Accuracy Sensitivity Specificity Precision F1 AUC Runtime min
CLIP Few-Shot 0.975 0.960 0.990 0.990 0.975 0.996 17.1
MobileNetV2-FS 0.725 0.510 0.940 0.895 0.650 0.583 1.3
AutoML 0.880 0.970 0.790 0.822 0.890 - - 203.0
- - indicates the value is not available/reported.